solve the equation
6-17/z -3/z^2=0

Answer by fractalier(6550)

Let us clear fractions by multiplying all by z^2, so from
6 - 17/z - 3/z^2 = 0 we get
6z^2 - 17z - 3 = 0 which we can factor into
(6z + 1)(z - 3) = 0 and solve, giving us
z = -1/6 and z = 3